OpenStudy (anonymous):

16*3^(4x-3) = 3*2^(8x-4)

OpenStudy (anonymous):

is this what the question look like \[16\times3^{(4x-3)}=3\times2^{(8x-4)}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, do you know the rules on indices?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

alright, we'll work through it step by step then first you need to solve 16x3 and 3x2

OpenStudy (anonymous):

49 and 6

OpenStudy (anonymous):

48

OpenStudy (anonymous):

yep so it should now look like \[48^{(4x-3)}=6^{(8x-4)}\]

OpenStudy (anonymous):

now the rules of indices when dealing in powers is that when the base are equal you can now eliminate the base and only work with the powers example \[3^{z}=3^{2}\] the because both side has a base number of 3 then z=2

OpenStudy (anonymous):

got that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so now we have to try make the base on both sides equal that is 48 and 6

OpenStudy (anonymous):

so we have to rewrite48 and 6 so that they have the same base?

OpenStudy (anonymous):

yeah, something like that. it might be a little had since both numbers cannot be squared

OpenStudy (anonymous):

wait a minute lets take a step back actually

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

see how you have 16x2 and 3x2 at first we have 2 as base in both

OpenStudy (anonymous):

where did you 16x2

OpenStudy (anonymous):

get

OpenStudy (anonymous):

i mean 16x3. sorry typing error

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so if we simplify 16 to \[2^{4}\] and 3x 2 stay s that way both base becomes \[3\times2^{4} and 3\times2\]

OpenStudy (anonymous):

so it is now \[3\times2^{4(4x-3)}=3\times2^{(8x-4)}\]

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

now we can do 4x-3 = 8x-4 ?

OpenStudy (anonymous):

yeah but remember the 4 so it should be 4(4x-3)=8x-4

OpenStudy (anonymous):

oh ok 16x-12 = 8x-4 so x = 1

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

absolutely..

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

so u can now solve for x

OpenStudy (anonymous):

your welcome